In excersise 33-41,Use the given information to find the given amount.
39. Original Price: $15
Discount Percent:20%
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

15-.2*15=15-3=$12 ANSWER FOR THE REDUCED PRICE.